You are here: Foswiki>PartnerObserving Web>ItalyQueue2012>ObsInaf20121210 (28 Feb 2013, EleanoraSani)Edit Attach

INAF LBC+MODS+LUCI queue observing, 2012 December 9-10

Observer: E. Sani, F. Cusano, M. Fumana
Telescope Operator: G. Becchetti
Support Astronomer: O. Kuhn
Instrument: LUCI

Summary

At the beginning of the twilight there are strong wind gusts. The Dome remains closed
Clear sky at the beginning of the observations with 69% humidity and wind speed 10m/s.
Poor to bad seeing during the night.
Data from luci.20121210.0001 to luci.20121210.0386.fits are calibration test performed in the afternoon.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Sozzetti_22 Wasp33 HKspec, 200_H+K 1h50m At the beginning of the occultation the dome is closed. Bad seeing. Problems with LUCI shield T.
Fritz_04 G190 dual-grating 1h40m Poor seeing between 1.5" and 2.0" but the targets are detected.

05.06 The T.O. opens the Dome

05.11 executeLuciscript.sh clean1.txt

05:33 LUCI reaches shield temperature limit.

05.36 executeLuciscript.sh spec_acq_science_22NEW.text luci.20121210.0387.fits-0393
  • We wait for the primary morrors to thermalize.

06:05 Acquisition done. Flushing detector to remove persistence from saturated target slit
  • Flush interupted after 20 exposures

06:10 executeLuciscript.sh spec_science_A_22.text
  • Filter wheel error. Run script again

06:15 executeLuciscript.sh spec_science_A_22.text luci.20121210.0394-0502
  • at the beginning of the integration the average seeing is 1.6

06.49 Flexure compensation problem. The problem is fixed without stopping the integration script.
  • The flexure compensation is again on luci.20121210.0449

07:20 Seeing very bad, guide seeing ~3

07:30 We are loosing guide. Script interrupted.
  • We miss about 4 minutes of integration

07:31 Run script again, guiding is not reliable luci.20121210.0503-0612
  • Seeing ~4"

08:17 Seeing ~2.7

08:47 No tracking, telescope loose the star, but preset not yet cancel.
  • Script interrupts with error: ERROR WITH OFFSET TELESCOPE COMMAND

08:50 Restart of temperature LUCI controller from Germany (this should prevent shield temperature warning)

08:54 We are waiting for the guiding star on the GSC monitor
  • Seeing 4.4"

09:01 Restarting of temperature controller seems to have no effect

09:25 We are not yet guiding and the occultation is over. We slew to LaFranca_A5

09:26 executeLuciscript.sh STDA5_1_J.acq
  • Filter wheel error

09:40 Shield temperature is raising again. No stable. We are check with the S.A. if we are able to still using LUCI or not

09:46 SWITCH TO MODS

10.31 We are authirized to use MODS
  • acqMods cxoj09.acq
  • modsAlign -l mods1r20121210.0002 mods1r.201210.0003
  • we are not able to able to see the target in the acquisition field even with a tipled exposure time. This is possibly due to the bad seeing.

10.55 We Slew to Fritz_04D
  • acqMods GG190.acq
  • modsAlign -l mods1r.20121210.0004.fits mods1r.20121210.0005.fits

11.00 execMods GG190.obs mods1r.20101210.0008 / 0009 mods1b.20121210.0002 / 0003
  • at the beginning of the scientific integration the average seeing is 1.5"

12:48 Script interrupted after 2 exposures because of incoming twilight

13.10 We are authorised to use LUCI and perform the K-band flat and lamps fro LaFranca_A11 in the K-band.
  • executeLuciscript.sh All_1_210_K.flat luci.20121210.0613-0652
  • executeLuciscript.sh All_1_210_K.arc luci.20121210.0653-0670

-- EleanoraSani - 10 Dec 2012
Topic revision: r5 - 28 Feb 2013, EleanoraSani
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback